### v2

# Variability study of classical supergiant X-ray binary 4U 1907+09 using NuSTAR

Abstract
We investigate the X-ray variability of the supergiant X-ray binary 4U 1907+09 using the new NuSTAR observation of 2024. The source had a relatively stable flux level during previous NuSTAR observations, but the flux varied significantly during the current one. The light curve exhibits dips (off-state) and flares (on-state). The phase-coherent timing analysis during the on-state yields a pulse period of $443.99(4)~{}\mathrm{s}$, showing the pulsar’s continued spin-down. The pulse profiles show an asymmetric double-peaked structure with a phase separation of 0.47 between the two peaks. A cyclotron resonance scattering feature (CRSF) is also detected at $\sim 17.6~{}\mathrm{keV}$, along with its harmonic at $\sim 38~{}\mathrm{keV}$, persisting across all flux states. Flux-resolved spectroscopy reveals that the CRSF remains constant despite a 25-fold change in flux. The spectral parameters like photon index and e-fold energy are out of phase with the pulse shape, whereas cutoff energy is in phase with the pulse shape. The source’s luminosity during the on-state is $2.85\times 10^{35}~{}\mathrm{erg~{}s^{-1}}$, consistent with a “pencil” beam radiation pattern expected at this flux level from a collisionless gas-mediated shock. These results offer further insights into the accretion dynamics and magnetic field geometry of this system.
keywords:
accretion , X-ray binaries , neutron star X-rays binaries , Accreting pulsar: 4U 1907+09 , data analysis
journal: arXiv
[…]
[label1]organization=Astrophysical Sciences Division, Bhabha Atomic Research Centre, city=Mumbai, postcode=400085, country=India
[…]
1 Introduction
Supergiant X-ray binaries (sgXBs) are binary systems that consist of a compact object and a massive O/B supergiant star. The accretion of stellar wind from the supergiant star onto a compact object is responsible for the X-ray emission in these systems. There are two types of X-ray binaries: classical supergiant high-mass binaries and supergiant fast X-ray transients (SFXTs) (Pradhan et al., 2018). 4U 1907+09 is a classical supergiant high-mass X-ray binary identified in the third Uhuru survey (Giacconi et al., 1971). The pulsar orbiting companion with an eccentricity $e\sim 0.28$ and an orbital period of $P\sim 8.3753$ d (in ’t Zand et al., 1998). Optical and infrared studies suggest an O8-O9 Ia type supergiant donor (Cox et al., 2005; Nespoli et al., 2008). The literature reports a source distance of 5 kpc (Cox et al., 2005; Nespoli et al., 2008). However, the distance to the source calculated by Gaia EDR3 is around 1.9 kpc (Bailer-Jones et al., 2021).
X-ray pulsation with a period of 437.5 s was initially reported by Makishima et al. (1984) utilizing Tenma observations. Early observations of the source revealed a relatively constant spin-down rate (Baykal et al., 2006); however, further observations revealed several torque reversals and spin-down rate variations (Fritz et al., 2006; Inam et al., 2009). INTEGRAL and RXTE measurements show short-term fluctuations in the pulse period compared to long-term changes in spin period rates, consistent with the random walk model (Şahiner et al., 2012). The most recent studies reported spin down of the pulsar (Varun et al., 2019; Tobrej et al., 2023).
4U 1907+09 is a variable X-ray source that exhibits irregular flaring and dipping activities. During the dips, in ’t Zand et al. (1997) did not detect X-ray pulsed emission from 4U 1907+09 using RXTE observations. The typical duration of the dips ranges from a few minutes to 1.5 hours. They suggested that cessation of accretion from the inhomogeneous wind of the companion star may be responsible for dips. Doroshenko et al. (2012) investigated the dipping activity using Suzaku observations and found the pulsations during the off state. The transition between off-state (dips) and on-state (normal flux state) was explained within the framework of “gated accretion”.
The X-ray spectra of 4U 1907+09 are characterized by a power law and an exponential cut-off Schwartz et al. (1980); Marshall and Ricketts (1980); Makishima et al. (1984); Cook and Page (1987); Chitnis et al. (1993); Roberts et al. (2001); Coburn et al. (2002); Baykal et al. (2006); Fritz et al. (2006); Varun et al. (2019); Tobrej et al. (2023). During the binary orbit, inhomogeneous accretion via intense stellar wind caused highly variable hydrogen column density ($n_{H}$) ranging from $1\times 10^{22}$ to $9\times 10^{22}$ cm-2 in ’t Zand et al. (1997). Moreover, the spectra showed a narrow spectral line at around 6.4 keV, which is the Fe $\mathrm{K}\alpha$ emission caused by the fluorescence of materials near the pulsar. Cyclotron-resonant scattering features (CRSFs) at $\sim 19$ keV with a second harmonic at $\sim 39$ keV have been seen using Ginga (Mihara, 1995) and BeppoSAX observations (Cusumano et al., 1998). This implies a surface magnetic field strength of $2.1\times 10^{12}$ G (Cusumano et al., 1998). The flaring activity observed by AstroSat indicates that the CRSF parameters have not been altered (Varun et al., 2019). The existence of cyclotron lines at $\sim 18$ keV and $\sim 38$ keV were recently verified by Tobrej et al. (2023) utilizing NuSTAR. Furthermore, an absorption line possibly due to Ni $\mathrm{K}\alpha$ and Ni $\mathrm{K}\beta$ at $\sim 8$ keV was also noticed.
In this work, we have estimated the spin rate and examined the flux-resolved spectral analysis of 4U 1907+09 using NuSTAR data observed on 20 November 2024. The manuscript is organized as follows: Section 2 describes observation and data reduction. Section 3 presents the timing and spectral data analysis results. Section 4 discusses the results. Section 5 summarises our findings.
2 Observation and data reduction
NuSTAR is a NASA space mission that observes the sky in the hard X-ray energy band (3-79 keV) of the electromagnetic spectrum (Harrison et al., 2013). It comprises two co-aligned grazing incidence telescopes with specially coated optics and detector units (FPMA and FPMB). In this work, we analyzed the NuSTAR data of 4U 1907+09 observed on MJD 60634.
HEASOFT version 6.33 and NuSTARDAS V2.1.2 with CALDBVER 20241126 were used to process and filter the NuSTAR event data. We selected a 90 arcsec radius region around the source centroid for the source and a 90 arcsec radius region away from the source for background estimation. The science products: spectra, light curves, and response files were extracted using the nuproducts script for both modules (FPMA and FPMB). We extracted the light curves with a time resolution of 0.1 s in the 3-79 keV energy range. The NuSTAR/FPMA light curve in the 3-79 keV energy range is shown in Figure 1. The source is largely variable during the observation. We divided the data into off-state and on-state. We extracted spectra and response files for the on and off states. We further divided the on-state into low and high luminosity and extracted the spectra and response files. The grppha task was used to rebin the source spectra for a minimum of 30 counts per bin.
Figure 1: NuSTAR light curve of 4U 1907+09 in 3-79 keV energy range for FPMA. The grey-shaded region shows the “off-state”. The blue and green-shaded region shows the “on-state” with low and high flux, respectively. (see section  3)
Figure 2: Pulse period history of 4U 1907+09. The pulse period measurements are take from Cook and Page (1987); in ’t Zand et al. (1998); Baykal et al. (2001, 2006); Mukerjee et al. (2001); Fritz et al. (2006); Inam et al. (2009); Şahiner et al. (2012); Varun et al. (2019); Tobrej et al. (2023). The red star corresponds to the spin period estimated in this work. (see section 3.1)
Figure 3: left panel: Pulse profile of 4U 1907+09 in 3-79 keV energy range using FPMA. Right panel: The pulse profile of 4U 1907+09 in high and low flux on-state. The green color corresponds to the low flux state and the blue color corresponds to the high flux state. (sec section 3.1)
We extracted the light curves during on state with a time resolution of 0.1 s in the following energy ranges: 3–10 keV, 10–15 keV, 15–20 keV, 20–30 keV, 30–50 keV, and 3–79 keV. We used barycorr tool to correct the photon arrival times to the barycenter of the solar system using the DE-430 ephemeris (Folkner et al., 2014).
3 Results
During this NuSTAR observation, the source was found to have different X-ray fluxes. We divided the data into off-state and on-state. The effective on-state and off-state exposure are 4.5 ks and 16.5 ks, respectively. Further, we divided the on state into high flux and low flux states. The effective exposure during high flux and low flux in the on state are 4.1 ks and 12.4 ks, respectively. The light curve and different states are shown in Figure 1. First, we performed temporal and spectral analyses on the averaged on-state.
3.1 Timing analysis
We examined the 3-79 keV light curve of averaged on-state for coherent pulsations from the source. We employed the xronos tool efsearch to estimate the best rotational period. The best period measured in the 3-79 keV energy range is 443.99(4)(MJD 60634.05). The uncertainty in the period was estimated using the bootstrap method by simulating 1000 light curves (Boldin et al., 2013). The calculated pulse period value with pulse period history of 4U 1907+09 is shown in Figure 2. The source was found to be in a spin-down phase. We employed the efold technique (Leahy et al., 1983) to create folded light curves at the best period in 3-79 keV, 3-10 keV, 10-15 keV, 15-20 keV, 20-30 keV and 30-50 keV. The left panel of Figure 3 shows the pulse profile in the 3-79 keV energy range. Figure 4 depicts the pulse shape in different energy ranges. There is a noticeable energy dependence in the pulse profile. A double-peaked feature is visible in the 3-79 keV energy range. This feature dominates in the 3-10 and 10-15 keV energy range. The peak at $\sim$ 0.25 starts to fade after the 15 keV. The pulse fraction (PF) is calculated using the expression,
$\mathrm{PF}=\frac{(I_{\mathrm{max}}-I_{\mathrm{min}})}{(I_{\mathrm{max}}+I_{\mathrm{min}})}$
Where $I_{\mathrm{max}}$ is the maximum normalized luminosity and $I_{\mathrm{min}}$ is the minimum normalized luminosity. The energy-dependent PF is shown in Figure 5. We checked the flux-dependent pulse shape as shown in the right panel of Figure 3. The pulse shape is similar in both low and high flux on-states.
Figure 4: Energy-dependent pulse profiles of 4U 1907+09 using FPMA. The energy range for the pulse profiles is specified inside the panels. (see section 3.1)
Figure 5: The energy dependence of pulse fraction. The dashed red line corresponds to the pulse fraction in the 3-79 keV energy range. (see section 3.1)
3.2 Spectral analysis
3.2.1 Averaged on state spectral analysis
We used XSPEC V12.14.0h (Arnaud, 1996) for spectral modeling of the data. A power-law model along with an energy cutoff and smoothing function, is used to represent the continuum spectra observed in X-ray pulsars. We fitted the data by model: const*tbabs(highecut*pow). The model const cross-calibrates instruments FPMA and FPMB. tbabs is Tuebingen-Boulder interstellar medium absorption model. pow is the photon power-law model. highecut is the high energy cutoff model given by,
$M(E)=\begin{array}[]{ll}exp\left[(E_{c}-E)/E_{f}\right]&E\geq E_{c}\\[5.69046pt] 1.0&E\leq E_{c}\end{array}$
where $E_{c}$ is cutoff energy (keV) and $E_{f}$ is e-folding energy (keV). The data does not effectively constrain the column density. So, we fixed the $n_{H}$ value at $1.5\times 10^{22}\;\mathrm{cm^{-2}}$ (Maitra and Paul, 2013; Varun et al., 2019). The $\chi^{2}/\mathrm{dof}$ comes out to be $\sim 2202/1119$. An iron line and absorption features around 18 keV and 35 keV due to cyclotron absorption are visible in the residual. So, we added model gauss for the Fe feature and a multiplicative model gabs. So, we added model gauss for the Fe feature and a multiplicative model gabs. The model becomes const*tbabs(gauss+highecut*pow)*gabs*gabs. The $\chi^{2}/\mathrm{dof}$ improved to $\sim 1277/1110$. The best fitting gives the $\Gamma\sim 1.12$ and $E_{\mathrm{cutoff}}\sim 11.8$ keV and $E_{\mathrm{fold}}\sim 12.8$ keV. The Gaussian line is fitted at 6.31 keV with an equivalent width equal to 46 eV, and the absorption lines came out at 17.6 keV and 38.02 keV. The flux came out to be $6.59_{-0.04}^{+0.04}\times 10^{-10}\;\mathrm{erg/cm^{2}/s}$.
To assess the importance of the absorption features in the spectra, we used the simftest script from xspec. It uses Monte Carlo simulations to produce simulated data sets from the original data, which are then used to estimate $\Delta\chi^{2}$ for gabs components. We conducted $10^{4}$ simulations for each gabs component. The $\Delta\chi^{2}$ obtained from simulated data sets is considerably low compared to the observed value for each gabs component. It confirms the significant detection of absorption features at $\sim 17.6$ keV and $\sim 38.02$ keV.
Then we fitted the data by model const*tbabs(gauss+compTT)*gabs*gabs. compTT (Titarchuk, 1994) is an analytic model describing the Comptonization of soft photons in hot plasma and is suitable for describing the continuum shape from X-ray Pulsars. The seed soft photon temperature, plasma temperature, and plasma optical depth are $\sim 0.52$ keV, $\sim 7.87$ keV, and $\sim 6.26$, respectively. The $\chi^{2}/\mathrm{dof}$ for this model is $\sim 1285/1110$ which is slightly higher than the previous model.
Finally, we employed the compmag (Farinelli et al., 2012) model to fit the continuum spectra. The compmag model describes the spectral formation in the accretion column onto the polar cap of a magnetized neutron star, taking into account both thermal and bulk Comptonization processes. Since the magnetic field of 4U 1907+09 is $2\times 10^{12}$ G, the compmag model is suitable for this system. The model becomes const*tbabs(gauss+bbodyrad+compmag)*gabs*gabs. During the fitting, we consider that the accretion velocity increases towards the NS (by fixing betaflag=1) with the index of velocity profile $\eta=0.5$ and terminal velocity at the NS surface $\beta_{0}=0.05$. We fixed the accretion column $r_{0}=0.25$ (in units of the NS Schwartzchild radius) and NS albedo A = 1. Since the norm of compmag is given by $R^{2}_{km}/D^{2}_{10}$, where $R_{km}$ is the radius of the accretion column in km and $D_{10}$ is the distance of the source in units of 10 kpc. Considering the radius of the accretion column is 1 km and the source distance is equal to 1.9 kpc. We fixed the norm at 28. The temperature of seed blackbody temperature ($kT_{bb}$) of compmag is tied with the blackbody temperature of bbodyrad. The remaining parameters of this spectral model were kept free. The $\chi^{2}/\mathrm{dof}$ for this model is $1236/1110$. The optical depth of the accretion column comes out to be $\sim 1.75$. Since the compmag model is applicable for $\tau>1$, we did not further use this model. Also, the low luminosity of the source disfavors the formation of the accretion column. The spectral parameters for different models used in this work are given in Table 1. The NuSTAR spectra with residuals of different models used in this work are shown in Figure 6.
Figure 6: Top: The NuSTAR spectra of 4U 1907+09. The green and red colour corresponds to FPMA and FPMB data, respectively. Second panel: residual of the continuum only model-const*tbabs(highecut*pow). The emission line at 6.5 keV and absorption features at 18 keV and 35 keV are visible. Third panel: residual of the model-const*tbabs(gauss+highecut*pow)gabs*gabs. Fourth panel: residual of the model-const*tbabs(gauss++compTT)gabs*gabs. Bottom panel: residual of the model-const*tbabs(gauss+bbodyrad+compmag)gabs*gabs. (see section 3.2)
3.2.2 Phase resolved Spectroscopy
We extracted the spectral files into eight phases: 0-0.125, 0.125-0.25, 0.25-0.375, 0.375-0.5, 0.5-0.625, 0.625-0.75, 0.75-875, and 0.875-1.0. We fitted the spectral data with the model: const*tbabs*(highecut*pow)gabs. The variation of different spectral parameters with phase is given in Figure 7. The top panel provides the 3.0-25.0 keV flux at various pulse phases, facilitating comparison with the pulse profile. The continuum fit parameters exhibit substantial fluctuation with the pulse phase. The cutoff energy and CRSF line center are correlated (in phase) with the pulse shape, whereas the e-fold energy and the photon index are out of phase with the pulse profile.
Figure 7: Phase resolved spectroscopy of 4U 1907+09. The flux is calculated in the 3-25 keV energy range. (see section 3.2.2)
3.2.3 Flux resolved spectral analysis
As shown in Figure 1, 4U 1907+09 shows the flux variation in the light curve. The spectral data for the off-state, low-flux on-state, and high-flux on-state is shown in Figure 8. We fitted the off-state and high-flux on-state data with the model const*tbabs(bbodyrad+highecut*pow)*gabs. The low flux on-state data required additional models gabs and gauss. We fitted the low flux on-state data with the model. const*tbabs(gauss+bbodyrad+highecut*pow)*
gabs*gabs
. The $\Gamma$ for off-state, high flux on-state, and low flux on-state come out to be $\sim 1.43$, $\sim 1.12$, and $\sim 1.08$, respectively. At various flux levels, the cyclotron line is found at $\sim 18$ keV. The line width of CRSF during high flux on-state and low flux on-state are $\sim 3.07$ and $\sim 2.5$, respectively. Compared to low flux on state, flux increased 3 times during high flux on state, and flux decreased 8 times during off-state. The e-folding energy during the high flux on-state and low flux on-state are $\sim 4.6$ and $\sim 6.27$, respectively. The spectral parameters for each data are given in Table 2.
Figure 8: The spectral data of 4U 1907+09 in off-state, low and high flux on-state. (see section 3.2.3)
Table 1: Best-fitting spectral parameters of 4U 1907+09 for different spectral models. The hydrogen column density ($N_{H}$) is fixed at $1.5\times 10^{22}\;\mathrm{cm^{-2}}$. Errors are reported at 90% confidence level.
Components Parameters M1 M2 M3
Const $C_{\mathrm{FPMB}}$ $1.003_{-0.006}^{+0.006}$ $1.004_{-0.006}^{+0.006}$ $1.004_{-0.005}^{+0.005}$
Powerlaw $\Gamma$ $1.12_{-0.10}^{+0.10}$
$\mathrm{Norm}$ $3.12_{-0.06}^{+0.06}\times 10^{-2}$
highecut $E_{\mathrm{cutoff}}$(keV) $11.8_{-0.3}^{+0.3}$
$E_{\mathrm{fold}}$ (keV) $12.8_{-1.0}^{+1.8}$
compTT $T0$ (keV) - $0.52_{-0.15}^{+0.07}$
$kT_{\mathrm{e}}$ (keV) - $7.87_{-1.38}^{+2.50}$
$\mathrm{\tau}$ $6.26_{-0.17}^{+0.27}$
norm (×$\times 10^{-2}$) - $1.39_{-0.08}^{+0.21}$
compmag $kT_{\mathrm{e}}$ (keV) - - $3.40_{-0.10}^{+0.15}$
$\mathrm{\tau}$ $1.75_{-0.14}^{+0.19}$
bbodyrad $kT_{\mathrm{bb}}$ (keV) - - $0.90_{-0.04}^{+0.03}$
norm - - $18.5_{-2.6}^{+5.0}$
gabs1 $E_{\mathrm{CRSF}}$(keV) $17.60_{-0.24}^{+0.12}$ $17.95_{-0.14}^{+0.16}$ $17.60_{-0.24}^{+0.30}$
$\sigma_{\mathrm{CRSF}}$ (keV) $2.23_{-0.36}^{+0.47}$ $2.88_{-0.13}^{+0.15}$ $2.68_{-0.27}^{+0.42}$
strength $1.78_{-0.36}^{+0.47}$ $4.47_{-0.36}^{+0.51}$ $3.06_{-0.52}^{+1.19}$
gabs2 $E_{\mathrm{CRSF}}$(keV) $38.02_{-0.82}^{+0.89}$ $39.55_{-1.29}^{+0.72}$ $34.93_{-0.66}^{+0.94}$
$\sigma_{\mathrm{CRSF}}$ (keV) $5.88_{-0.80}^{+0.98}$ $10.00_{-1.10}^{+1.33}$ $4.04_{-0.82}^{+1.67}$
strength $28.44_{-5.97}^{+8.88}$ $84.41_{-25.15}^{+8.89}$ $11.20_{-2.94}^{+7.12}$
gauss $E_{\mathrm{Fe}}$ (keV) $6.31_{-0.06}^{+0.05}$ $6.29_{-0.07}^{+0.06}$ $6.18_{-0.07}^{+0.11}$
$\sigma_{\mathrm{Fe}}$ (keV) $0.11_{-0.11}^{+0.12}$ $0.24_{-0.11}^{+0.13}$ $0.64_{-0.15}^{+0.16}$
norm (×$\times 10^{-4}$) $1.82_{-0.45}^{+0.52}$ $2.78_{-0.62}^{+0.73}$ $7.91_{-0.22}^{+0.25}$
EQW (eV) 46 70 207
$\chi^{2}$/dof 1277/1110 1285/1110 1236/1110
M1: const*tbabs(gauss+highecut*powerlaw)*gabs*gabs
M2: const*tbabs(gauss+compTT)*gabs*gabs
M3: const*tbabs(gauss+bbodyrad+compmag)*gabs*gabs
Table 2: Best-fitting spectral parameters for off-state, low flux on-state and high flux on-state. The hydrogen column density ($N_{H}$) is fixed at $1.5\times 10^{22}\;\mathrm{cm^{-2}}$. The unit of flux is (×$(\times 10^{-10})\;\mathrm{erg\;cm^{-2}\;s^{-1}}$. Errors are reported at 90% confidence level.
Components Parameters off-state High flux on-state Low-flux on-state
Const $C_{\mathrm{FPMB}}$ $0.977_{-0.041}^{+0.043}$ $0.996_{-0.008}^{+0.008}$ $1.009_{-0.009}^{+0.009}$
Powerlaw $\Gamma$ $1.43_{-0.07}^{+0.07}$ $1.12_{-0.02}^{+0.02}$ $1.08_{-0.02}^{+0.02}$
$\mathrm{Norm}$ $4.34_{-0.53}^{+0.61}\times 10^{-3}$ $6.59_{-0.20}^{+0.21}\times 10^{-2}$ $1.84_{-0.05}^{+0.06}\times 10^{-2}$
highecut $E_{\mathrm{cutoff}}$(keV) $22.9_{-2.5}^{+1.1}$ $19.3_{-0.78}^{+0.51}$ $18.26_{-0.50}^{+0.41}$
$E_{\mathrm{fold}}$ (keV) $4.6_{-1.5}^{+2.8}$ $4.6_{-0.25}^{+0.37}$ $6.27_{-0.42}^{+0.56}$
gabs1 $E_{\mathrm{CRSF}}$(keV) $18.56_{-0.61}^{+0.64}$ $18.31_{-0.43}^{+0.31}$ $18.09_{-0.23}^{+0.24}$
$\sigma_{\mathrm{CRSF}}$ (keV) $2.20_{-0.44}^{+0.53}$ $3.07_{-0.24}^{+0.19}$ $2.50_{-0.13}^{+0.15}$
strength $3.43_{-0.99}^{+1.10}$ $6.64_{-0.84}^{+0.67}$ $5.05_{-0.43}^{+0.45}$
gabs2 $E_{\mathrm{CRSF}}$(keV) - - $33.69_{-0.83}^{+0.95}$
$\sigma_{\mathrm{CRSF}}$ (keV) - - $2.99_{-0.76}^{+1.28}$
strength - - $5.68_{-2.00}^{+3.62}$
gauss $E_{\mathrm{Fe}}$ (keV) - - $6.26_{-0.08}^{+0.07}$
$\sigma_{\mathrm{Fe}}$ (keV) - - $0.16_{-0.13}^{+0.11}$
norm (×$\times 10^{-4}$) - - $1.54_{-0.41}^{+0.46}$
[…]
$\chi^{2}$/dof 205/181 1045/929 1047/940
4 Discussion
4U 1907+09 is a classical supergiant X-ray binary known to show the dips and flares in its light curve. In this work, we analyzed the NuSTAR data of 4U 1907+09 observed during November 2024. The spin down/spin up and torque reversals of the pulsar were documented in several studies (Cook and Page, 1987; in ’t Zand et al., 1998; Baykal et al., 2001, 2006; Mukerjee et al., 2001; Fritz et al., 2006; Inam et al., 2009; Şahiner et al., 2012; Varun et al., 2019; Tobrej et al., 2023). In the most recent study, Tobrej et al. (2023) found that the 4U 1907+09 has been spinning down. In this work, we estimated the pulse period of 4U 1907+09 to be $\sim 443.99(4)$ s. This indicates that the source is still spinning down. The pulse profile is one of the indicators of the emission region from the neutron star. A double-peaked (at $\sim 0.26$ and $0.73$) asymmetric pulse shape was observed in 4U 1907+09, as shown in the left panel of Figure 3. A similar pulse shape was also reported for the 4U 1907+09 in the previous studies (Maitra and Paul, 2013; Varun et al., 2019; Tobrej et al., 2023). The observed asymmetric pulse shape is most likely caused by a magnetic dipole offset from the rotational axis of the pulsar (Parmar et al., 1989; Leahy, 1991; Riffert et al., 1993; Bulik et al., 1995). Another plausible explanation for the asymmetric shape of pulse profiles is an asymmetric accretion stream (Basko and Sunyaev, 1976; Wang and Welter, 1981; Miller, 1996). As energy increases, the pulse shape changes significantly. The peak at 0.26 begins to fade at 15 keV. The PF followed an increasing behaviour with energy.
The NuSTAR spectrum was fitted with three different models: power law with highecut, compTT, and compmag. Table 1 contains the corresponding fit parameters. The iron line and CRSF with its harmonic were found at $\sim 6.31$ keV, $\sim 17.6$ keV, and $\sim 38.02$ keV, respectively. After incorporating gravitational redshift, the CRSF corresponds to a magnetic field of $\sim 2\times 10^{12}$ G. The average flux of the source during the on-state in the 3–50 keV energy range was found to be $\sim 6.59\times 10^{-10}\mathrm{erg\;cm^{-2}\;s^{-1}}$. Assuming a distance of 1.9 kpc, the estimated luminosity of the source is $\sim 2.85\times 10^{35}\mathrm{erg\;s^{-1}}$. The luminosity of the source is similar to that observed by NuSTAR in 2018 (Tobrej et al., 2023). Using the formalism in Becker et al. (2012), we estimate the critical luminosity $L_{\mathrm{crit}}=1.5\times 10^{37}\;\left(\frac{B}{10^{12}\;\mathrm{G}}\right)^{16/15}\;\mathrm{erg\;s^{-1}}$ to be $3\times 10^{37}\mathrm{erg\;s^{-1}}$. The Coulomb braking luminosity $L_{\mathrm{coul}}=1.2\times 10^{37}\;\left(\frac{B}{10^{12}\;\mathrm{G}}\right)^{1/3}\;\mathrm{erg\;s^{-1}}$ to be $1.5\times 10^{37}\mathrm{erg\;s^{-1}}$ (Becker et al., 2012; Langer and Rappaport, 1982). The observed luminosity of 4U 1907+09 is much lower than the $L_{\mathrm{crit}}$ and $L_{\mathrm{coul}}$. As a result, radiation from a gas-mediated shock without collision is anticipated. This leads to a ”pencil” beam, where the radiation is emitted mostly along the magnetic field lines.
Numerous studies examine the dips and flares in the light curve of 4U 1907+09 (in ’t Zand et al., 1997; Doroshenko et al., 2012; Şahiner et al., 2012; Shakura et al., 2012; Varun et al., 2019). Apart from 4U 1907+09, off-states or dips have been reported in the Vela X-1 (Kreykenbohm et al., 2008; Doroshenko et al., 2011) and GX 301-2 (Göğüş et al., 2011). Previous studies have ruled out the likelihood of dips generated by a dense, clumpy wind obscuring the neutron star (Fürst et al., 2011; Şahiner et al., 2012). Shakura et al. (2012) suggested that the onset of the off-state in these systems is due to the transition from the Compton cooling-dominated to the radiative cooling-dominated regime. During this observation, the source showed the dip and flare in the light curve. During the dip, the flux of the source decreased 8 times the flux during the low flux on-state. The flux during high flux on-state increased 3 times during low flux on-state. As the flux varied during the on-state, the pulse profile did not differ much and followed a similar trend (as shown in the right panel of Figure 3). It indicates that the emission process did not change during the flare activity. The spectral modelling of 4U 1907+09 revealed that the $E_{\mathrm{CRSF}}$ is unaltered at $\sim 18$ keV in both the high flux (flare) on-state and the off-state. It suggests that the accretion column may not be formed. The e-folding energy during the high flux on-state is lower than the low flux high state. It shows that the high flux on-state is softer than the low flux on-state. The line width of the CRSF during the high flux on-state is higher than the low flux on-state.
5 Summary
The results and findings of this work are summarized as follows:
  • 1.
    4U 1907+09 exhibiting normal flux, flare, and dip (off-state) in the light curve and providing insights into the variability of the source.
  • 2.
    With an estimated pulse period of $443.99(4)$ s, the source was determined to be spin down.
  • 3.
    During the on-state, the typical spectral and temporal characteristics of 4U 1907+09 are detected, including a double-peaked pulse profile, energy dependence of pulse profile, CRSF, and dips in the light curve.
  • 4.
    The CRSF at $\sim 18$ keV was found in both on and off states. The width of CRSF correlated positively with the flux.
  • 5.
    The spin phase influenced the spectral properties. Cutoff energy is in phase with the pulse shape, but photon index and e-fold energy are out of phase.
Acknowledgments
We have utilized the archived NuSTAR data provided by the High Energy Astrophysics Science Archive Research Center (HEASARC) online service maintained by the Goddard Space Flight Center. For this work, we have made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). RK acknowledges the Department of Astronomy and Astrophysics, Tata Institute of Fundamental Research, Mumbai, India for providing research facilities.